OpenStudy (anonymous):

Which equation represents the line that contains the point (-2, 9) and is perpendicular to the line y = 2x + 6? A. y = -(1/2)x – 7 B. y = -(1/2)x + 8 C. y = -(1/2)x + 11 D. y = 2x + 13

Nnesha (nnesha):

i already told you about perpendicular lines

OpenStudy (anonymous):

did i post this before? andi dont remember

Nnesha (nnesha):

ohh well this is same question

OpenStudy (anonymous):

ok whats the formula agian

OpenStudy (anonymous):

@Mesopretty

Nnesha (nnesha):

see your previous post now you should remember the formula

OpenStudy (anonymous):

i cant veiw it

Nnesha (nnesha):

why not ??

OpenStudy (anonymous):

idk cant find it

OpenStudy (anonymous):

Because this person only wants the answer?

OpenStudy (anonymous):

noo

Nnesha (nnesha):

okay this is like same question now you should remember the concept

OpenStudy (anonymous):

then stop being lazy and look at the old post.

OpenStudy (anonymous):

go away ryan

OpenStudy (anonymous):

is it y-y=m(x-x

Nnesha (nnesha):

try to be independent :) here is a formula y =mx+b

OpenStudy (anonymous):

ok thank you

Nnesha (nnesha):

yeah let me now what u get :) take your time

OpenStudy (anonymous):

whats b?

Nnesha (nnesha):

you have to find b

Nnesha (nnesha):

what is your slope of new equation ??

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

-m-2

Nnesha (nnesha):

nope i told you definition of perpendicular reciprocal 2 <---- flip it and change sign

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

idk

Nnesha (nnesha):

what you don't know ??

Nnesha (nnesha):

don't know how to flip 2/1 and change sign if its negative then change to positive and if is positive change to negative

OpenStudy (anonymous):

nevermind ill ask my teacher

Nnesha (nnesha):

that's good you really have to take notice when you ask question next time you should have notebook or pencil with you so then you can do same kinda question bec you don't want to look over previous post

Nnesha (nnesha):

notes**